let y= 2x^4 - 12x^2 + 2
a. find all point of inflection. be sure to write your answer in form (x,y)
b. list the intervals on which the graph is concave upward and concave downward.
Answer #1

2 I=207 – 12 x²+2 a = 8x²240 Jy2 = 242² 24 = 24(2-1) (2+1) @ paint of inflection are givenby day=0 > 24 (2-1) (2+1)=0 dar - 2